# Fuzzy Dark Matter Constraints from the Hubble Frontier Fields

### v1

@@ Abstract
In fuzzy dark matter (FDM) cosmologies, the dark matter consists of ultralight bosons ($m\lesssim 10^{-20}$ eV). The astrophysically large de Broglie wavelengths of such particles hinder the formation of low-mass dark matter halos. This implies a testable prediction: a corresponding suppression in the faint-end of the ultraviolet luminosity function (UVLF) of galaxies. Notably, recent estimates of the faint-end UVLF at $z\sim 5-9$ in the Hubble Frontier Fields, behind foreground lensing clusters, probe up to five magnitudes fainter than typical (“blank-field”) regions. These measurements thus far disfavor prominent turnovers in the UVLF at low luminosity, implying bounds on FDM. We fit a semi-empirical model to these and blank-field UVLF data, including the FDM particle mass as a free parameter. This fit excludes cases where the dark matter is entirely a boson of mass $m<1.5\times 10^{-21}$ eV (with $2\sigma$ confidence). We also present a less stringent bound deriving solely from the requirement that the total observed abundance of galaxies, integrated over all luminosities, must not exceed the total halo abundance in FDM. This more model-agnostic bound disfavors $m<5\times 10^{-22}$ eV ($2\sigma$). We forecast that future UVLF measurements from JWST lensing fields may probe masses several times larger than these bounds, although we demonstrate this is subject to theoretical uncertainties in modeling the FDM halo mass function.
@@ 1 Introduction
While there is much evidence the majority of matter in the universe is dark (i.e. does not interact appreciably with electromagnetic fields) (Bertone & Tait, 2018; Planck Collaboration et al., 2020; Arbey & Mahmoudi, 2021), its precise character remains elusive (Bertone & Tait, 2018; Cirelli et al., 2024). In this work we assess the possibility that the dark matter consists of fuzzy dark matter (FDM) particles (Hu et al., 2000), ultralight bosons with particle masses of $m\sim 10^{-22}-10^{-20}$ eV). [^1] This is the range of interest for our present analysis although the term may describe several decades in mass around this range. FDM is sometimes alternatively referred to as wave, axion-like, or Bose-Einstein condensate dark matter, as it is a subclass of these categories. The cosmological deBroglie wavelengths in FDM leave detectable signatures in the large-scale structure of the universe (see e.g. Marsh 2016a; Niemeyer 2020; Hui 2021; Ferreira 2021; O’Hare 2024 for reviews). We contrast this with conventional cold dark matter (CDM): a collisionless, non-relativistic species without wave-like effects on astrophysical scales.
@@
Early work (e.g. Hu et al. 2000, see Lee 2018 for a brief history) often considered FDM as a resolution of the so-called “small-scale” problems of CDM (e.g., the missing satellite, cuspy halo, and too-big-to-fail problems, see Del Popolo & Le Delliou (2017); Bullock & Boylan-Kolchin (2017) for reviews). While it is possible that these issues are resolved solely by the impact of baryonic feedback in CDM, it is important to recognize that CDM is less-sharply tested on small spatial scales. Alternative possibilities such as FDM make distinctive small-scale predictions and thereby provide a useful foil for CDM models. It may be that FDM is what is realized in nature. Indeed, FDM is well-motivated from particle physics considerations: ultralight scalar particles with the correct relic density to make up the dark matter (Marsh, 2016a; Hui et al., 2017; Hui, 2021) arise naturally in string theory (Arvanitaki et al., 2010) or owing to symmetry-breaking processes in the early universe, e.g. via the misalignment mechanism analogous to that of the proposed QCD axion (Peccei & Quinn, 1977; Wilczek, 1978; Weinberg, 1978). There are additional well-motivated alternatives to CDM besides FDM, including warm dark matter (WDM) in which free-streaming leads to a suppression of small scale fluctuations. Although our focus in this work is on FDM, we will briefly consider the implications of our results for WDM as well.
@@
The ultraviolet luminosity function (UVLF) describes the number density of galaxies as a function of their luminosity. The form of the UVLF depends on the underlying abundance of dark matter halos and how galaxy luminosity scales with host halo mass (Yang et al., 2003; Cooray & Milosavljević, 2005; Bouwens et al., 2015; Schive et al., 2016; Sipple & Lidz, 2024). Therefore it is possible to use the UVLF to test theories, like FDM, which predict small dark matter halos to be rare. Analyses using the gravitational lensing of galaxies in the Hubble Frontier Fields (HFF) (Coe et al., 2015; Lotz et al., 2017) have probed the UVLF at $z\leq 9$ at up to five magnitudes fainter (Livermore et al., 2017; Atek et al., 2018; Bhatawdekar et al., 2019; Bouwens et al., 2022b) than in unmagnified regions (known as the field or blank-field regions) (Bouwens et al., 2021). The faintest galaxies in the lensed samples reside in halos that are roughly an order of magnitude smaller in mass than any in the field (Sipple & Lidz, 2024), yet the UVLF data do not show an obvious faint-end suppression. As we will see, these measurements thereby tighten constraints on the FDM particle mass compared to previous UVLF studies (Schive et al., 2016; Corasaniti et al., 2017; Ni et al., 2019; Winch et al., 2024) making it competitive with similar bounds from observations of the Lyman-$\alpha$ forest (Iršič et al., 2017; Kobayashi et al., 2017; Armengaud et al., 2017; Rogers & Peiris, 2021) and Milky Way satellite galaxies (Schutz, 2020; Nadler et al., 2021; Banik et al., 2021). The consistency across multiple independent probes helps establish the robustness of the FDM constraints.
@@ 2.1 The Halo Mass Function in FDM
If the dark matter is primarily FDM its large scale evolution matches CDM, but structure formation near and below its de Broglie wavelength is opposed by the uncertainty principle. The balance between this effect and gravity sets a relevant Jeans scale, $k_{J}$, where there is significant deviation from CDM for $k\gtrsim k_{J}$ (Hu et al., 2000). In linear theory this co-moving Jeans wavenumber scales as $k_{J}\propto(1+z)^{-1/4}$ during matter domination (Hu et al., 2000), i.e. FDM suppression effects apply to smaller co-moving scales over time. Consequently, it is common to approximate FDM as evolving like CDM with a suppression in the initial linear power spectrum (at matter-radiation equality), ignoring subsequent dynamical effects which would modify the growth of the smaller scale modes. In most of this work we employ fitting formulas for the halo mass function derived from FDM simulations that use this approximation, as full FDM simulations in large cosmological volumes require a currently prohibitive dynamic range in spatial scale. In Section 5 we return to this point and consider the sensitivity of our results to this approximation.
@@
The halo mass function describes the number density of halos per mass interval. As structure grows from the initial density perturbations, the formation of halos near and below the half-mode mass scale $M_{1/2}=(4\pi\bar{\rho}_{m}/3)(\pi/k_{1/2})^{3}=5.6\times 10^{10}m_{22}^{-4/3}M_{\odot}$ will be suppressed. The FDM dynamical effects mentioned above, including complex non-linear wave effects, may also influence the shape and redshift evolution of the HMF (see e.g. Hui et al. 2017; Ferreira 2021). Current studies suggest that dynamical effects become important on scales where structure is already heavily suppressed (Zhang et al., 2018; Nori et al., 2019; Li et al., 2019; May & Springel, 2023), as alluded to above, and so they are often neglected in FDM HMF models.
@@
with $M_{0}=1.6\times 10^{10}m_{22}^{-4/3}M_{\odot}\approx 0.3M_{1/2}$. This matches CDM for $M>>M_{0}$, but drops to $\sim 80\%/20\%/0.5\%$ the CDM value for $M=10M_{0},M_{0},0.1M_{0}$. The halo suppression mass scale shrinks with increasing particle mass as $M_{0}\propto m_{22}^{-4/3}$. Qualitatively, these effects can be seen as a consequence of the uncertainty principle. As a larger particle mass implies a smaller de Broglie wavelength ($\lambda_{\rm dB}\propto\frac{1}{m}$) it is possible to confine the dark matter within smaller regions. In the context of the HMF, heavier FDM particles are more difficult to distinguish from CDM. In this parametrization CDM corresponds to the $m_{22}\rightarrow\infty$ limit.
We compare the CDM and FDM mass functions at $z=8$ in Figure 1 for the cases of $m_{22}=1,3,10,$ and $30$. The top panel shows the mass functions themselves while the bottom gives the ratio between each FDM model and CDM. The gray shaded regions illustrate the approximate mass scales probed by different UVLF measurement sets assuming the best-fit luminosity-halo mass relation of Sipple & Lidz (2024) (or, as we find equivalently, this work (Section 3)). The shaded regions, from dark to light, show: the field (unlensed) HST (B21) ($\gtrsim 3\times 10^{10}M_{\odot}$), lensed HST (B22) ($\gtrsim 2\times 10^{9}M_{\odot}$), and forecasted lensed JWST (Section 7) (Jaacks et al., 2019; Shen et al., 2023) ($\gtrsim 4\times 10^{8}M_{\odot}$). Observational probes which investigate the HMF at lower mass scales will be better able to distinguish between FDM scenarios and CDM. Equation 1 gives a redshift-independent ratio between the HMF in FDM and CDM. However, plausible alternative FDM halo mass function models predict more pronounced differences with CDM at higher redshift (see Section 5).
@@ 2.2.1 The Conditional Luminosity Function
where the five parameters, $L_{0}$, $M_{1}$, $p$, $q$, and $r$, are calibrated using UVLF measurements. The double power-law form reflects the expectation that feedback effects suppress star formation (and hence UV luminosities) most strongly in low-mass and high-mass halos (see Silk (2011) for a brief review). We discuss the impact of uncertainties in the role of feedback effects in Section 4.
The conditional distribution connecting the UVLF and HMF, $\phi_{c}(L|M,z)$, is taken to be a lognormal with scatter $\sigma_{\rm LN}=0.37$ and median $\mu_{\rm LN}=L_{c}(M,z)$. [^3] We fix the value of the scatter following previous treatments (Sipple & Lidz, 2024; Schive et al., 2016; Bouwens et al., 2015, 2008) as it is generally not well-constrained. Assuming each dark matter halo above some limiting mass $M_{\rm min}$ will, on average, host one luminous galaxy (supported by low satellite fractions at $z\sim 6$ (Harikane et al., 2022)) and taking a duty cycle of unity, the relation to the UVLF $\phi(L,z)$ [^4] In units of absolute magnitude, $M_{\rm UV}=51.6-2.5\log_{10}\left(L/[\text{erg}\text{s}^{-1}\text{Hz}^{-1}]\right)$ (Oke, 1974). Under change of variables: $\phi(L)\text{d}L=\phi(M_{\rm UV})\text{d}M_{\rm UV}$. is:
@@
Explicitly, the vector of parameters of our model fit is defined as ${\bm{\theta}}=\left(p,q,r,M_{\rm UV,0},{\rm log_{10}}(M_{1}/{\rm M}_{\odot}),1/m_{22}\right)$. Our priors on these parameters are: ([0.8, 3.5], [0.8, 2.5], [0, 2.5], [-25, -20], [10, 14], [0, 2]). We choose to sample in inverse $m_{22}$ to allow the CDM limit $m_{22}\rightarrow\infty$ to be explicitly considered in the prior volume. See Section 4.3 for additional discussion regarding the FDM mass prior. Likelihoods are determined following Sipple & Lidz (2024), where the likelihood of a model UVLF value at a given magnitude is Gaussian where symmetric errors are quoted, half-Gaussian when only upper limits are quoted, or follows the prescription of Equations 13-16 of Barlow (2004) for asymmetric errors.
@@ 3 Results
Table 1: Best-fit UVLF model parameters and 68% confidence intervals determined from MCMC sampling under different modelling assumptions. The first row displays the main results of our fiducial model, described in Sections 2 and 3. The next two rows show results when restricted to CDM-only scenarios or to not use the lensed UVLF measurements from B22 respectively. The following three rows are modified models described in Section 4: adding additional parameters to Equation 2 (Section 4.2) or with alternate priors for $m_{22}$ (Section 4.3) of either flat in $\log_{10}(m_{22})$ or based on the Jeffreys prior. The final two rows use the alternative FDM HMF parametrizations of Kulkarni & Ostriker (2021) (Section 5.1) or Marsh & Silk (2013); Marsh (2016b) (Section 5.2).
@@
Figure 3: Top: 1D marginal histogram of FDM mass in our fiducial model from MCMC sampling. $1/2/3\sigma$ limits are shown as broken vertical lines. Models of $m_{22}\gtrsim 100$, including CDM, are indistinguishable from one another using the current data and this region is prior-dominated. Bottom: 2D marginal distributions of $1/m_{22}$ with the other five parameters. Blue lines and rectangles show best-fit parameter values from scipy.optimize. There is little covariance with the other parameters, whose distributions differ little from the fit assuming CDM (see Table 1 and SL).
@@
There is almost no covariance between $1/m_{22}$ and the other parameters, with the largest correlation coefficient in magnitude being $\sim-0.16$ with the faint-end slope parameter $p$ [^6] Correlations between the five parameters of Equation 2 are similar to Sipple & Lidz (2024) (see their Figure 2 and Section 4.3) That is, at least in the context of a double power-law model, the FDM bound is not strongly subject to uncertainties in the relationship between UV luminosity and host halo mass. In this model, compensating for the sharp faint-end suppression caused by FDM necessarily comes at the expense of fitting the other measurements less well. In the following two sections we consider the robustness of these conclusions to a range of alternative assumptions, such as allowing additional freedom in the star formation efficiency of small halos or alternate models of the FDM HMF.

### v2

@@ Abstract
In fuzzy dark matter (FDM) cosmologies, the dark matter consists of ultralight bosons ($m\lesssim 10^{-20}$ eV). The astrophysically large de Broglie wavelengths of such particles hinder the formation of low-mass dark matter halos. This implies a testable prediction: a corresponding suppression in the faint-end of the ultraviolet luminosity function (UVLF) of galaxies. Notably, recent estimates of the faint-end UVLF at $z\sim 5-9$ in the Hubble Frontier Fields, behind foreground lensing clusters, probe up to five magnitudes fainter than typical (“blank-field") regions. These measurements thus far disfavor prominent turnovers in the UVLF at low luminosity, implying bounds on FDM. We fit a semi-empirical model to these and blank-field UVLF data, including the FDM particle mass as a free parameter. This fit excludes cases where the dark matter is entirely a boson of mass $m<1.5\times 10^{-21}$ eV (with $2\sigma$ confidence). We also present a less stringent bound deriving solely from the requirement that the total observed abundance of galaxies, integrated over all luminosities, must not exceed the total halo abundance in FDM. This more model-agnostic bound disfavors $m<5\times 10^{-22}$ eV ($2\sigma$). We forecast that future UVLF measurements from JWST lensing fields may probe masses several times larger than these bounds, although we demonstrate this is subject to theoretical uncertainties in modeling the FDM halo mass function.
keywords:
galaxies: luminosity function – galaxies: high-redshift – cosmology: theory – dark matter
pubyear: 2025pagerange: Fuzzy Dark Matter Constraints from the Hubble Frontier FieldsFuzzy Dark Matter Constraints from the Hubble Frontier Fields
@@ 1 Introduction
While there is much evidence the majority of matter in the universe is dark (i.e. does not interact appreciably with electromagnetic fields) (Bertone & Hooper, 2018; Planck Collaboration et al., 2020; Arbey & Mahmoudi, 2021), its precise character remains elusive (Bertone & Tait, 2018; Cirelli et al., 2024). In this work we assess the possibility that the dark matter consists of fuzzy dark matter (FDM) particles (Hu et al., 2000), ultralight bosons with particle masses of $m\sim 10^{-22}-10^{-20}$ eV). [^1] This is the range of interest for our present analysis although the term may describe several decades in mass around this range. FDM is sometimes alternatively referred to as wave, axion-like, or Bose-Einstein condensate dark matter, as it is a subclass of these categories. The cosmological de Broglie wavelengths in FDM leave detectable signatures in the large-scale structure of the universe (see e.g. Marsh 2016b; Schwabe 2018; Niemeyer 2020; Hui 2021; Ferreira 2021; O’Hare 2024 for reviews). We contrast this with conventional cold dark matter (CDM): a collisionless, non-relativistic species without wave-like effects on astrophysical scales.
@@
Early work (e.g. Hu et al. 2000, see Lee 2018 for a brief history) often considered FDM as a resolution of the so-called “small-scale” problems of CDM (e.g., the missing satellite, cuspy halo, and too-big-to-fail problems, see Del Popolo & Le Delliou (2017); Bullock & Boylan-Kolchin (2017) for reviews). While it is possible that that these issues owe partly to observational incompleteness (Kim et al., 2018) and/or may be resolved through including the impact of baryonic feedback in CDM (Sawala et al., 2016), it is important to recognize that CDM is less-sharply tested on small spatial scales. Alternative possibilities such as FDM make distinctive small-scale predictions and thereby provide a useful foil for CDM models. It may be that FDM is what is realized in nature. Indeed, FDM is well-motivated from particle physics considerations: ultralight scalar particles with the correct relic density to make up the dark matter (Marsh, 2016b; Hui et al., 2017; Hui, 2021) arise naturally in string theory (Arvanitaki et al., 2010) or owing to symmetry-breaking processes in the early universe, e.g. via the misalignment mechanism analogous to that of the proposed QCD axion (Peccei & Quinn, 1977; Wilczek, 1978; Weinberg, 1978). There are additional well-motivated alternatives to CDM besides FDM, including warm dark matter (WDM) in which free-streaming leads to a suppression of small scale fluctuations. Although our focus in this work is on FDM, we will briefly consider the implications of our results for WDM as well.
@@
The ultraviolet luminosity function (UVLF) describes the number density of galaxies as a function of their luminosity. The form of the UVLF depends on the underlying abundance of dark matter halos and how galaxy luminosity scales with host halo mass (Yang et al., 2003; Cooray & Milosavljević, 2005; Bouwens et al., 2015; Schive et al., 2016; Sipple & Lidz, 2024). Therefore it is possible to use the UVLF to test theories, like FDM, which predict small dark matter halos to be rare. Analyses using the gravitational lensing of galaxies in the Hubble Frontier Fields (HFF) (Coe et al., 2015; Lotz et al., 2017) have probed the UVLF at $z\leq 9$ at up to five magnitudes fainter (Livermore et al., 2017; Ishigaki et al., 2018; Atek et al., 2018; Bhatawdekar et al., 2019; Bouwens et al., 2022b) than in unmagnified regions (known as the field or blank-field regions) (Bouwens et al., 2021).
Lensed UVLF measurements are inherently challenging since magnification factor and survey volume estimates require modeling cluster mass distributions. Indeed, if unaccounted systematic errors remain in the UVLF measurements at low luminosity, this could influence our constraints. In this work we assume the binned UVLF measurements of Bouwens et al. (2022b) and refer the reader to that work and its companion, Bouwens et al. (2022a), for a more detailed discussion of potential uncertainties in determining the UVLF behind lensing clusters and progress in mitigating these effects. For instance, they marginalize over uncertainties in the cluster mass distribution model and reassuringly find consistency between the faint-end slope estimates from the field and behind the cluster lenses. Conservatively, we also include separate bounds using only the Bouwens et al. (2021) field measurements.
The faintest galaxies in the lensed samples reside in halos that are roughly an order of magnitude smaller in mass than any in the field (Sipple & Lidz, 2024), yet the UVLF data do not show an obvious faint-end suppression. As we will see, these measurements thereby tighten constraints on the FDM particle mass compared to previous UVLF studies (Schive et al., 2016; Corasaniti et al., 2017; Menci et al., 2017; Ni et al., 2019; Lapi et al., 2022; Winch et al., 2024) making it competitive with similar bounds from observations of the Lyman-$\alpha$ forest (Iršič et al., 2017; Kobayashi et al., 2017; Armengaud et al., 2017; Rogers & Peiris, 2021) and Milky Way satellite galaxies (Schutz, 2020; Nadler et al., 2021; Banik et al., 2021). The consistency across multiple independent probes helps establish the robustness of the FDM constraints.
@@ 2.1 The Halo Mass Function in FDM
If the dark matter is primarily FDM its large scale evolution matches CDM, but structure formation near and below its de Broglie wavelength is opposed by the uncertainty principle. The balance between this effect and gravity sets a relevant Jeans scale, $k_{J}$, where there is significant deviation from CDM for $k\gtrsim k_{J}$ (Khlopov et al., 1985; Hu et al., 2000). In linear theory this co-moving Jeans wavenumber scales as $k_{J}\propto(1+z)^{-1/4}$ during matter domination (Hu et al., 2000), i.e. FDM suppression effects apply to smaller co-moving scales over time. Consequently, it is common to approximate FDM as evolving like CDM with a suppression in the initial linear power spectrum (at matter-radiation equality), ignoring subsequent dynamical effects which would modify the growth of the smaller scale modes. In most of this work we employ fitting formulas for the halo mass function derived from FDM simulations that use this approximation, as full FDM simulations in large cosmological volumes require a currently prohibitive dynamic range in spatial scale. In Section 5 we return to this point and consider the sensitivity of our results to this approximation.
@@
The halo mass function describes the number density of halos per mass interval. As structure grows from the initial density perturbations, the formation of halos near and below the half-mode mass scale $M_{1/2}=(4\pi\bar{\rho}_{m}/3)(\pi/k_{1/2})^{3}=5.6\times 10^{10}m_{22}^{-4/3}M_{\odot}$ will be suppressed. The FDM dynamical effects mentioned above, including complex non-linear wave effects, may also influence the shape and redshift evolution of the HMF (see e.g. Hui et al. 2017; Ferreira 2021). Current studies suggest that dynamical effects become important on scales where structure is already heavily suppressed (Zhang et al., 2018b; Nori et al., 2019; Li et al., 2019; May & Springel, 2023), as alluded to above, and so they are often neglected in FDM HMF models.
@@
with $M_{0}=1.6\times 10^{10}m_{22}^{-4/3}M_{\odot}\approx 0.3M_{1/2}$. This matches CDM for $M>>M_{0}$, but drops to $\sim 80\%/20\%/0.5\%$ the CDM value for $M=10M_{0},M_{0},0.1M_{0}$. The halo suppression mass scale shrinks with increasing particle mass as $M_{0}\propto m_{22}^{-4/3}$. Qualitatively, these effects can be seen as a consequence of the uncertainty principle. As a larger particle mass implies a smaller de Broglie wavelength $\lambda_{\rm dB}\propto 1/m$) it is possible to confine the dark matter within smaller regions. In the context of the HMF, heavier FDM particles are more difficult to distinguish from CDM. In this parametrization CDM corresponds to the $m_{22}\rightarrow\infty$ limit.
We compare the CDM and FDM mass functions at $z=8$ in Figure 1 for the cases of $m_{22}=1,3,10,$ and $30$. The top panel shows the mass functions themselves while the bottom gives the ratio between each FDM model and CDM. The gray shaded regions illustrate the approximate mass scales probed by different UVLF measurement sets assuming the best-fit luminosity-halo mass relation of Sipple & Lidz (2024) (or, as we find equivalently, this work (Section 3)). The shaded regions, from dark to light, show: the field (unlensed) HST (B21) ($\gtrsim 3\times 10^{10}M_{\odot}$), lensed HST (B22) ($\gtrsim 2\times 10^{9}M_{\odot}$), and forecasted lensed JWST (Section 7) (Jaacks et al., 2019; Shen et al., 2024) ($\gtrsim 4\times 10^{8}M_{\odot}$). Observational probes which investigate the HMF at lower mass scales will be better able to distinguish between FDM scenarios and CDM. Equation 1 gives a redshift-independent ratio between the HMF in FDM and CDM. However, plausible alternative FDM halo mass function models predict more pronounced differences with CDM at higher redshift (see Section 5).
@@ 2.2.1 The Conditional Luminosity Function
where the five parameters, $L_{0}$, $M_{1}$, $p$, $q$, and $r$, are calibrated using UVLF measurements. The double power-law form reflects the expectation that feedback effects suppress star formation (and hence UV luminosities) most strongly in low-mass and high-mass halos (see Silk (2011) for a brief review). [^3] Other approaches, such as GALLUMI (Sabti et al., 2022), instead model the ratio of stellar mass to halo mass (or the ratio of their time derivatives) as a double power-law in halo mass. Such models should be comparable to the conditional luminosity function approach used in this paper, as UV luminosity and star formation rate are well-correlated and proportional to one another (Madau & Dickinson, 2014) (see e.g. Sipple & Lidz 2024 for a discussion on the connection of the conditional luminosity function to other physical quantities). We explore the impact of uncertainties in the role of feedback effects in Section 4.
The conditional distribution connecting the UVLF and HMF, $\phi_{c}(L|M,z)$, is taken to be a lognormal with scatter $\sigma_{\rm LN}=0.37$ and median $\mu_{\rm LN}=L_{c}(M,z)$. [^4] We fix the value of the scatter following previous treatments (Sipple & Lidz, 2024; Schive et al., 2016; Bouwens et al., 2015, 2008) as it is generally not well-constrained. Assuming each dark matter halo above some limiting mass $M_{\rm min}$ will, on average, host one luminous galaxy (supported by low satellite fractions at $z\sim 6$ (Harikane et al., 2022)) and taking a duty cycle of unity, the relation to the UVLF $\phi(L,z)$ is:
@@
[…]
@@ 3 Results
Table 1: Best-fit UVLF model parameters and 68% confidence intervals determined from MCMC sampling under different modelling assumptions. The first row displays the main results of our fiducial model, described in Sections 2 and 3. The next two rows show results when restricted to CDM-only scenarios or to not use the lensed UVLF measurements from B22 respectively. The following three rows are modified models described in Section 4: adding additional parameters to Equation 2 (Section 4.2) or with alternate priors for $m_{22}$ (Section 4.3) of either flat in $\log_{10}(m_{22})$ or based on the Jeffreys prior. The final two rows use the alternative FDM HMF parametrizations of Kulkarni & Ostriker (2022) (Section 5.1) or Marsh & Silk (2014); Marsh (2016a) (Section 5.2).
@@
Figure 3: Top: 1D marginal histogram of FDM mass in our fiducial model from MCMC sampling. The $1/2/3\sigma$ limits are shown as broken vertical lines. Models with $m_{22}\gtrsim 100$, including CDM, are indistinguishable from one another using the current data and this region is prior-dominated. Bottom: 2D marginal distributions of $1/m_{22}$ with the other five parameters. Blue lines and rectangles show best-fit parameter values from scipy.optimize. There is little covariance with the other parameters, whose distributions differ little from the fit assuming CDM (see Table 1 and SL).
@@
There is almost no covariance between $1/m_{22}$ and the other parameters, with the largest correlation coefficient in magnitude being $\sim-0.16$ with the faint-end slope parameter $p$ [^7] Correlations between the five parameters of Equation 2 are similar to Sipple & Lidz (2024) (see their Figure 2 and Section 4.3). That is, at least in the context of a double power-law model, the FDM bound is not strongly subject to uncertainties in the relationship between UV luminosity and host halo mass. To illustrate, consider the extreme case of the $m_{22}=2$ curve in Figure 2. In this model a single power law index covers a wide range in luminosity and so fitting galaxies in the range $-16\lesssim M_{\rm UV}\lesssim-13$ would necessarily come at the expense of fitting $-20\lesssim M_{\rm UV}\lesssim-16$ less well. Additionally, while the relative suppression of the HMF in FDM asymptotes to a power law at low mass (see Equation 1), its intermediate effect will not be perfectly compensated for by a corresponding power law enhancement in the luminosity-halo mass relation. This further motivates an exploration into our model assumptions, which we undertake in the rest of the paper.
In the following two sections we consider the robustness of these conclusions to a range of alternative assumptions, such as allowing additional freedom in the star formation efficiency of small halos or alternate models of the FDM HMF.